a web designer is trying to put 4 characters on a page. the space occupied by character 2 takes 5 units more than character 1. character 3 takes 7 units less than character 2. character 4 takes the same amount of space as character 2. i have to find the total space occupied by all the characters in an algebraic expression
Answer by stanbon(48568)

a web designer is trying to put 4 characters on a page. the space occupied by character 2 takes 5 units more than character 1. character 3 takes 7 units less than character 2. character 4 takes the same amount of space as character 2. i have to find the total space occupied by all the characters in an algebraic expression
-----------
Let space for character 1 be "x".
2 space = x+5
3 space = x+5-7=x-2
4 space = x+5
--------------
Total space = x + x+5 + x-2 + x+5 = 4x+8
=============
Cheers,
Stan H.